# Unveiling Multiple Descents in Unsupervised Autoencoders

### v2

@@ Results.
Ensuring the robustness of the findings across various model initializations and enhancing their reliability, all figures combine several results of different random seeds. The bolded curves in each figure represent the average across these results, and the transparent curve around each bolded curve represents the ±$\pm 1$ standard error from the mean.
@@ Appendix B SNR Calculations
${\boldsymbol{z}}-d\times 1$ vector. Represents a vector in a latent space of size $d$.
${\boldsymbol{H}}-D\times d$ matrix. Represents a random matrix to project ${\boldsymbol{z}}$ from a $d$ dimensional space into a higher-dimensional space ($D>d$).
$\boldsymbol{\epsilon}-D\times 1$ vector. Represents the noise added to a vector with $D$ dimensions.
@@
$\text{SNR}^{2}=\frac{E[\|\beta{\boldsymbol{H}}{\boldsymbol{z}}\|_{2}^{2}]}{E[\|\boldsymbol{\epsilon}\|_{2}^{2}]}=\frac{E[\beta^{2}{\boldsymbol{z}}^{\top}{\boldsymbol{H}}^{\top}{\boldsymbol{H}}{\boldsymbol{z}}]}{E[\boldsymbol{\epsilon}^{\top}\boldsymbol{\epsilon}]}=\frac{\beta^{2}E_{\boldsymbol{z}}[E_{{\boldsymbol{H}}|{\boldsymbol{z}}}[{\boldsymbol{z}}^{\top}{\boldsymbol{H}}^{\top}{\boldsymbol{H}}{\boldsymbol{z}}|{\boldsymbol{z}}]]}{E\left[\sum_{i=1}^{D}\boldsymbol{\epsilon}_{i}^{2}\right]}\underset{(1)}{=}$ (3)
$\frac{\beta^{2}E_{\boldsymbol{z}}[{\boldsymbol{z}}^{\top}E_{{\boldsymbol{H}}|{\boldsymbol{z}}}[{\boldsymbol{H}}^{\top}{\boldsymbol{H}}]{\boldsymbol{z}}]}{D}\underset{(2)}{=}\frac{\beta^{2}E_{\boldsymbol{z}}[{\boldsymbol{z}}^{\top}D\cdot\mathbb{I}_{d\times d}{\boldsymbol{z}}]}{D}=\frac{\beta^{2}\cdot D\cdot E_{\boldsymbol{z}}[{\boldsymbol{z}}^{\top}{\boldsymbol{z}}]}{D}=\beta^{2}E\left[\sum_{i=1}^{d}{\boldsymbol{z}}_{i}^{2}\right]\underset{(1)}{=}\beta^{2}\cdot d.$
@@
(1) Given a vector ${\boldsymbol{a}}\sim\mathcal{N}(0,\mathbb{I}_{D})$ of $D$ i.i.d. samples, $E\left[\sum_{i=1}^{D}{\boldsymbol{a}}_{i}^{2}\right]=\sum_{i=1}^{D}E[{\boldsymbol{a}}_{i}^{2}]=\sum_{i=1}^{D}1=D$.
(2) Given a matrix ${\boldsymbol{M}}\sim\mathcal{N}(0,\mathbb{I}_{D})$ of size $D\times D$ where all entries are i.i.d., then
$E[{\boldsymbol{M}}^{\top}{\boldsymbol{M}}]=E\begin{bmatrix}{\boldsymbol{M}}_{1,1}^{2}+\dots+{\boldsymbol{M}}_{D,1}^{2}&\dots&{\boldsymbol{M}}_{1,1}{\boldsymbol{M}}_{1,D}+\dots+{\boldsymbol{M}}_{D,1}{\boldsymbol{M}}_{D,D}\\ \vdots&\ddots&\vdots\\ {\boldsymbol{M}}_{1,1}{\boldsymbol{M}}_{1,D}+\dots+{\boldsymbol{M}}_{D,1}{\boldsymbol{M}}_{D,D}&\dots&{\boldsymbol{M}}_{1,D}^{2}+\dots+{\boldsymbol{M}}_{D,D}^{2}\end{bmatrix}=\begin{bmatrix}D&\dots&0\\ \vdots&\ddots&\vdots\\ 0&\dots&D\end{bmatrix}=D\cdot\mathbb{I}_{D\times D}$.
For the scenario of feature noise, each train sample has only $D\cdot p$ noisy features, meaning the noise vector contains values for only $D\cdot p$ entries. Consequently, $\beta$ is determined by $\sqrt{\frac{p}{d}}\cdot\text{SNR}$. For practitioners who want to explore the scenario involving domain shift, where the source and target are noisy, note that the matrix responsible for projecting ${\boldsymbol{z}}_{test}$ into a higher-dimensional space is denoted as ${\boldsymbol{H}}^{{}^{\prime\prime}}={\boldsymbol{H}}+s\cdot{\boldsymbol{H}}^{{}^{\prime}}$ where ${\boldsymbol{H}}^{{}^{\prime}}$ is sampled from a standard normal distribution $\mathcal{N}(0,\mathbb{I})$ and both ${\boldsymbol{H}}$ and ${\boldsymbol{H}}^{{}^{\prime}}$ are i.i.d. Consequently, ${\boldsymbol{H}}_{ij}^{{}^{\prime\prime}}\sim\mathcal{N}(0,1+s^{2})$. Substituting ${\boldsymbol{H}}$ with ${\boldsymbol{H}}^{{}^{\prime\prime}}$ in equation (3), we find that $E_{{\boldsymbol{H}}^{{}^{\prime\prime}}|{\boldsymbol{z}}}[{\boldsymbol{H}}^{{}^{\prime\prime}\top}{\boldsymbol{H}}^{{}^{\prime\prime}}]=D\cdot(1+s^{2})\cdot\mathbb{I}_{d}$, leading to $\text{SNR}^{2}=(1+s^{2})\cdot\beta^{2}\cdot d$, therefore $\beta=\frac{\text{SNR}}{\sqrt{(s^{2}+1)d}}$. In other words, since the covariance matrix of ${\boldsymbol{H}}^{{}^{\prime\prime}}$ is $(1+s^{2})\mathbb{I}$, we need to make sure we first normalize the matrix by $\sqrt{1+s^{2}}$ to maintain the identity covariance matrix.
For other datasets, such as the single-cell RNA dataset, we normalize each sample ${\boldsymbol{x}}$ by its norm $\|{\boldsymbol{x}}\|$, and similarly normalize each noise vector $\boldsymbol{\epsilon}$, yielding: $\hat{{\boldsymbol{x}}}=\dfrac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}$ and $\hat{\boldsymbol{\epsilon}}=\dfrac{\boldsymbol{\epsilon}}{\|\boldsymbol{\epsilon}\|}$. This ensures that the ratio $\dfrac{\hat{{\boldsymbol{x}}}}{\hat{\boldsymbol{\epsilon}}}$ equals 1. By employing equation (2), we attain the intended linear SNR factor $\beta$, and then scale down $\hat{\boldsymbol{\epsilon}}$ by $\beta$, yielding $\hat{\boldsymbol{\epsilon}}_{scaled}=\dfrac{\hat{\boldsymbol{\epsilon}}}{\beta}$. This guarantees that the linear SNR is $\dfrac{\hat{{\boldsymbol{x}}}}{\hat{\boldsymbol{\epsilon}}_{scaled}}=\beta$.
@@
Let ${\boldsymbol{X}}^{D\times N}=({\boldsymbol{x}}_{1},{\boldsymbol{x}}_{2},\ldots,{\boldsymbol{x}}_{N})$ represent the training set, with each column ${\boldsymbol{x}}_{i}$ corresponding to a data point according to the formulation in equation (1). Assume that the number of samples $N$ is large enough such that the rank of the data matrix ${\boldsymbol{X}}$ is almost surely equal to the number of features $D$. This assumption is reasonable due to the noise added to many samples.
Theorem $\mathbf{4.1}$ [No double descent in linear Autoencoders] Let $\Phi_{a}(\cdot;\boldsymbol{\theta}),\,\Phi_{b}(\cdot;\boldsymbol{\theta}):\mathbb{R}^{D}\to\mathbb{R}^{D}$ be two linear AEs, each with $L\geq 2$ layers and architectures defined by $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ (refer to section 4 for definition), with bottleneck sizes $m_{a}$ and $m_{b}$, where $m_{a}<m_{b}\leq D$. Let $\boldsymbol{\theta}_{a}^{\text{opt}}$ and $\boldsymbol{\theta}_{b}^{\text{opt}}$ be the optimal parameters under assumption C.1. Let $d$ denote the dimension of the support of $\mathcal{P}_{\tilde{{\boldsymbol{x}}}}$ (the dimension of the span of the samples $\tilde{{\boldsymbol{x}}}\sim\mathcal{P}_{\tilde{{\boldsymbol{x}}}}$). Then, the following holds,
$\mathcal{G}\left(\mathcal{D}_{b};\boldsymbol{\theta_{b}}^{\text{opt }},\mathcal{P}_{\tilde{{\boldsymbol{x}}}}\right)\leq\mathcal{G}\left(\mathcal{D}_{a};\boldsymbol{\theta_{a}}^{\text{opt }},\mathcal{P}_{\tilde{{\boldsymbol{x}}}}\right),$
@@
For simplicity of notation, we will omit the attributions $a$ and $b$ unless explicitly needed. Without the loss of generality, we denote the training set by ${\boldsymbol{X}}^{D\times N}=({\boldsymbol{x}}_{1},{\boldsymbol{x}}_{2},\ldots,{\boldsymbol{x}}_{N})$ where the matrix column ${\boldsymbol{x}}_{i}$ are w.r.t formulation (1). Recall that $\text{rank}({\boldsymbol{X}})=D,$ with probability $1$ under assumption C.1. Let $\boldsymbol{\Sigma}\in\mathbb{R}^{D\times D}$ represent the covariance matrix, defined as $\boldsymbol{\Sigma}={\boldsymbol{X}}{\boldsymbol{X}}^{\top}$. Let its singular value decomposition (SVD) be expressed as $\boldsymbol{\Sigma}={\boldsymbol{U}}\boldsymbol{\Lambda}{\boldsymbol{U}}^{\top}$, and similarly $\boldsymbol{\Sigma}_{\tilde{{\boldsymbol{x}}}}=\tilde{{\boldsymbol{U}}}\boldsymbol{\Omega}\tilde{{\boldsymbol{U}}}^{\top}$. Note that the eigenvectors of $\boldsymbol{\Sigma}$, which are the columns of ${\boldsymbol{U}}$, are arranged in descending order corresponding to their associated eigenvalues in $\boldsymbol{\Lambda}$. By Lemma C.2, since $\text{rank}({\boldsymbol{X}})=D\geq m$, then for the optimal solution, $\boldsymbol{\theta}^{opt}=\text{argmin}_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};{\boldsymbol{X}}),$ we have that $\operatorname{rank}\left({\boldsymbol{W}}^{opt}_{L-1}\cdots{\boldsymbol{W}}^{opt}_{2}\right)=m,$ means that the rank of the inner layers multiplication is exactly the size of the bottleneck. By Theorem 3.2, proof Case II in (Kawaguchi, 2016), all local minima of the considered optimization problem are global minima, and the expression of a global minimum of the training optimization, namely
$\boldsymbol{\theta}^{opt}=\text{argmin}_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};{\boldsymbol{X}})=\text{argmin}_{\boldsymbol{\theta}}\frac{1}{N}\|\Phi({\boldsymbol{X}};\boldsymbol{\theta})-{\boldsymbol{X}}\|_{F}^{2},$
@@
$\boldsymbol{\theta}^{opt}={\boldsymbol{U}}_{m}{\boldsymbol{U}}_{m}^{\top}{\boldsymbol{X}}{\boldsymbol{X}}^{\top}\left({\boldsymbol{X}}{\boldsymbol{X}}^{\top}\right)^{-},$
where ${\boldsymbol{U}}_{m},$ is the matrix containing the first $m$ eigenvectors of the full eigenbasis ${\boldsymbol{U}}$ of ${\boldsymbol{X}}{\boldsymbol{X}}^{\top},$ namely ${\boldsymbol{U}}_{m}=\left[{\boldsymbol{u}}_{1},\ldots,{\boldsymbol{u}}_{m}\right]$ (those corresponding to the $m$ largest eigenvalues), and the operator $(\cdot)^{-}$ stands for the pesudo inverse. Since $\boldsymbol{\Sigma}={\boldsymbol{X}}{\boldsymbol{X}}^{\top}$ is invertible ($\text{rank}({\boldsymbol{X}})=D$), we have
$\boldsymbol{\theta}^{opt}={\boldsymbol{U}}_{m}{\boldsymbol{U}}_{m}^{\top},$
@@
$\Phi({\boldsymbol{X}};\boldsymbol{\theta})={\boldsymbol{U}}_{m}{\boldsymbol{U}}_{m}^{\top}{\boldsymbol{X}}.$
Due to the misalignment, ${\boldsymbol{U}}$ and $\tilde{{\boldsymbol{U}}}$ are related by a unitary matrix ${{\boldsymbol{R}}}$, such that,
${\boldsymbol{U}}={{\boldsymbol{R}}}\tilde{{\boldsymbol{U}}}.$
@@
$\mathcal{G}(\mathcal{D};\boldsymbol{\theta}^{opt},\mathcal{P}_{\tilde{{\boldsymbol{x}}}})$ $=\mathbb{E}\left[\|\Phi(\tilde{{\boldsymbol{x}}};\boldsymbol{\theta}^{opt})-\tilde{{\boldsymbol{x}}}\|_{2}^{2}\right]$
$=\mathbb{E}\left[\|{\boldsymbol{U}}_{m}{\boldsymbol{U}}_{m}^{\top}\tilde{{\boldsymbol{x}}}-\tilde{{\boldsymbol{x}}}\|_{2}^{2}\right]$
$=\mathbb{E}\left[\|\right(\mathbb{I}-{\boldsymbol{U}}_{m}{\boldsymbol{U}}_{m}^{\top}\left)\tilde{{\boldsymbol{x}}}\|_{2}^{2}\right]$
$=\mathbb{E}\left[\tilde{{\boldsymbol{x}}}^{\top}\left(\mathbb{I}-{{\boldsymbol{R}}}\tilde{{\boldsymbol{U}}}_{m}\tilde{{\boldsymbol{U}}}_{m}^{\top}{{\boldsymbol{R}}}^{\top}\right)\tilde{{\boldsymbol{x}}}\right]$
$=\mathbb{E}\left[\|\tilde{{\boldsymbol{x}}}\|_{2}^{2}\right]-\mathbb{E}\left[\left\|\tilde{{\boldsymbol{U}}}_{m}^{\top}{{\boldsymbol{R}}}^{\top}\tilde{{\boldsymbol{x}}}\right\|_{2}^{2}\right]$
$=\mathrm{Tr}(\boldsymbol{\Sigma}_{\tilde{{\boldsymbol{x}}}})-\mathbb{E}\left[\left\|\tilde{{\boldsymbol{U}}}_{m}^{\top}{{\boldsymbol{R}}}^{\top}\tilde{{\boldsymbol{x}}}\right\|_{2}^{2}\right],$ (4)
where $\boldsymbol{\Sigma}_{\tilde{{\boldsymbol{x}}}}:=\mathbb{E}\left[\tilde{{\boldsymbol{x}}}\tilde{{\boldsymbol{x}}}^{\top}\right]$. Now, since there exists at least $d$ different indexes $j\in[D]$ for which ${\boldsymbol{u}}_{j}^{\top}{{\boldsymbol{R}}}\neq 0$ (otherwise the latent dimension of the test-set $\tilde{{\boldsymbol{x}}}$ would be strictly lower than $d$ in contradiction), then for $m_{a}+1<d$ and $m_{b}+1<d$, $\mathbb{E}\left[\left\|\tilde{{\boldsymbol{U}}}_{m_{a}}^{\top}{{\boldsymbol{R}}}^{\top}\tilde{{\boldsymbol{x}}}\right\|_{2}^{2}\right],\quad\mathbb{E}\left[\left\|\tilde{{\boldsymbol{U}}}_{m_{b}}^{\top}{{\boldsymbol{R}}}^{\top}\tilde{{\boldsymbol{x}}}\right\|_{2}^{2}\right]>0.$ Hance, since the larger is $m$ in (4) the lower is $\mathcal{G},$ we have
$\mathcal{G}\left(\mathcal{D}_{b};\boldsymbol{\theta_{b}}^{\text{opt }},\mathcal{P}_{\tilde{{\boldsymbol{x}}}}\right)\leq\mathcal{G}\left(\mathcal{D}_{a};\boldsymbol{\theta_{a}}^{\text{opt }},\mathcal{P}_{\tilde{{\boldsymbol{x}}}}\right),$
@@
$\mathcal{G}\left(\mathcal{D}_{b};\boldsymbol{\theta_{b}}^{\text{opt }},\mathcal{P}_{\tilde{{\boldsymbol{x}}}}\right)=\mathcal{G}\left(\mathcal{D}_{a};\boldsymbol{\theta_{a}}^{\text{opt }},\mathcal{P}_{\tilde{{\boldsymbol{x}}}}\right)=0.$
@@
Consider the pre-trained linear autoencoder, $\Phi(\cdot;\boldsymbol{\theta}^{\text{opt}}):\mathbb{R}^{D}\to\mathbb{R}^{D}$, of $L\geq 2$ layers, where $\boldsymbol{\theta}^{\text{opt}}\equiv\left\{{\boldsymbol{W}}_{1}^{d_{1}\times D},\ldots,{\boldsymbol{W}}_{L-1}^{d_{L-1}\times d_{L-2}},{\boldsymbol{W}}_{L}^{D\times d_{L-1}}\right\}$, for which its bottleneck size holds $m(\leq D)$, and a training set ${\boldsymbol{X}},$ of any dimension satisfies $\text{rank}\left({\boldsymbol{X}}\right)\geq m,$ then $\operatorname{rank}\left({\boldsymbol{W}}_{L-1}\cdots{\boldsymbol{W}}_{2}\right)=m.$

### v3

@@ Results.
Ensuring the robustness of the findings across various model initializations and enhancing their reliability, all figures combine several results of different random seeds. The bolded curves in each figure represent the average across these results, and the transparent curve around each bolded curve represents the $\pm 1$ standard error from the mean.
@@ Appendix B SNR Calculations
${\bm{z}}-d\times 1$ vector. Represents a vector in a latent space of size $d$.
${\bm{H}}-D\times d$ matrix. Represents a random matrix to project $\bm{z}$} from a $d$ dimensional space into a higher-dimensional space ($D>d$).
$\bm{\epsilon}-D\times 1$ vector. Represents the noise added to a vector with $D$ dimensions.
@@
$\text{SNR}^{2}=\frac{E[\|\beta{\bm{H}}{\bm{z}}\|_{2}^{2}]}{E[\|\bm{\epsilon}\|_{2}^{2}]}=\frac{E[\beta^{2}{\bm{z}}^{\top}{\bm{H}}^{\top}{\bm{H}}{\bm{z}}]}{E[\bm{\epsilon}^{\top}\bm{\epsilon}]}=\frac{\beta^{2}E_{\bm{z}}[E_{{\bm{H}}|{\bm{z}}}[{\bm{z}}^{\top}{\bm{H}}^{\top}{\bm{H}}{\bm{z}}|{\bm{z}}]]}{E\left[\sum_{i=1}^{D}\bm{\epsilon}_{i}^{2}\right]}\underset{(1)}{=}$ (3)
$\frac{\beta^{2}E_{\bm{z}}[{\bm{z}}^{\top}E_{{\bm{H}}|{\bm{z}}}[{\bm{H}}^{\top}{\bm{H}}]{\bm{z}}]}{D}\underset{(2)}{=}\frac{\beta^{2}E_{\bm{z}}[{\bm{z}}^{\top}D\cdot\mathbb{I}_{d\times d}{\bm{z}}]}{D}=\frac{\beta^{2}\cdot D\cdot E_{\bm{z}}[{\bm{z}}^{\top}{\bm{z}}]}{D}=\beta^{2}E\left[\sum_{i=1}^{d}{\bm{z}}_{i}^{2}\right]\underset{(1)}{=}\beta^{2}\cdot d.$
@@
(1) Given a vector ${\bm{a}}\sim\mathcal{N}(0,\mathbb{I}_{D})$ of $D$ i.i.d. samples, $E\left[\sum_{i=1}^{D}{\bm{a}}_{i}^{2}\right]=\sum_{i=1}^{D}E[{\bm{a}}_{i}^{2}]=\sum_{i=1}^{D}1=D$.
(2) Given a matrix ${\bm{M}}\sim\mathcal{N}(0,\mathbb{I}_{D})$ of size $D\times D$ where all entries are i.i.d., then
$E[{\bm{M}}^{\top}{\bm{M}}]=E\begin{bmatrix}{\bm{M}}_{1,1}^{2}+\dots+{\bm{M}}_{D,1}^{2}&\dots&{\bm{M}}_{1,1}{\bm{M}}_{1,D}+\dots+{\bm{M}}_{D,1}{\bm{M}}_{D,D}\\ \vdots&\ddots&\vdots\\ {\bm{M}}_{1,1}{\bm{M}}_{1,D}+\dots+{\bm{M}}_{D,1}{\bm{M}}_{D,D}&\dots&{\bm{M}}_{1,D}^{2}+\dots+{\bm{M}}_{D,D}^{2}\end{bmatrix}=\begin{bmatrix}D&\dots&0\\ \vdots&\ddots&\vdots\\ 0&\dots&D\end{bmatrix}=D\cdot\mathbb{I}_{D\times D}$.
For the scenario of feature noise, each train sample has only $D\cdot p$ noisy features, meaning the noise vector contains values for only $D\cdot p$ entries. Consequently, $\beta$ is determined by $\sqrt{\frac{p}{d}}\cdot\text{SNR}$. For practitioners who want to explore the scenario involving domain shift, where the source and target are noisy, note that the matrix responsible for projecting ${\bm{z}}_{test}$ into a higher-dimensional space is denoted as ${\bm{H}}^{{}^{\prime\prime}}={\bm{H}}+s\cdot{\bm{H}}^{{}^{\prime}}$ where $\bm{H}$$^{{}^{\prime}}$ is sampled from a standard normal distribution $\mathcal{N}(0,\mathbb{I})$ and both $\bm{H}$} and $\bm{H}$$^{{}^{\prime}}$ are i.i.d. Consequently, ${\bm{H}}_{ij}^{{}^{\prime\prime}}\sim\mathcal{N}(0,1+s^{2})$. Substituting $\bm{H}$} with $\bm{H}$$^{{}^{\prime\prime}}$ in equation equation 3, we find that $E_{{\bm{H}}^{{}^{\prime\prime}}|{\bm{z}}}[{\bm{H}}^{{}^{\prime\prime}\top}{\bm{H}}^{{}^{\prime\prime}}]=D\cdot(1+s^{2})\cdot\mathbb{I}_{d}$, leading to $\text{SNR}^{2}=(1+s^{2})\cdot\beta^{2}\cdot d$, therefore $\beta=\frac{\text{SNR}}{\sqrt{(s^{2}+1)d}}$. In other words, since the covariance matrix of $\bm{H}$$^{{}^{\prime\prime}}$ is $(1+s^{2})\mathbb{I}$, we need to make sure we first normalize the matrix by $\sqrt{1+s^{2}}$ to maintain the identity covariance matrix.
For other datasets, such as the single-cell RNA dataset, we normalize each sample $\bm{x}$} by its norm $\|{\bm{x}}\|$, and similarly normalize each noise vector $\bm{\epsilon}$, yielding: $\hat{{\bm{x}}}=\dfrac{{\bm{x}}}{\|{\bm{x}}\|}$ and $\hat{\bm{\epsilon}}=\dfrac{\bm{\epsilon}}{\|\bm{\epsilon}\|}$. This ensures that the ratio $\dfrac{\hat{{\bm{x}}}}{\hat{\bm{\epsilon}}}$ equals 1. By employing equation equation 2, we attain the intended linear SNR factor $\beta$, and then scale down $\hat{\bm{\epsilon}}$ by $\beta$, yielding $\hat{\bm{\epsilon}}_{scaled}=\dfrac{\hat{\bm{\epsilon}}}{\beta}$. This guarantees that the linear SNR is $\dfrac{\hat{{\bm{x}}}}{\hat{\bm{\epsilon}}_{scaled}}=\beta$.
@@
Let ${\bm{X}}^{D\times N}=({\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N})$ represent the training set, with each column ${\bm{x}}_{i}$ corresponding to a data point according to the formulation in equation (1). Assume that the number of samples $N$ is large enough such that the rank of the data matrix $\bm{X}$} is almost surely equal to the number of features $D$. This assumption is reasonable due to the noise added to many samples.
Theorem $\mathbf{4.1}$ [No double descent in linear AEs] Let $\Phi_{a}(\cdot;\bm{\theta}),\,\Phi_{b}(\cdot;\bm{\theta}):\mathbb{R}^{D}\to\mathbb{R}^{D}$ be two linear AEs, each with $L\geq 2$ layers and architectures defined by $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ (refer to section 5 for definition), with bottleneck sizes $m_{a}$ and $m_{b}$, where $m_{a}<m_{b}\leq D$. Let $\bm{\theta}_{a}^{\text{opt}}$ and $\bm{\theta}_{b}^{\text{opt}}$ be the optimal parameters under assumption C.1. Let $d$ denote the dimension of the support of $\mathcal{P}_{\tilde{{\bm{x}}}}$ (the dimension of the span of the samples $\tilde{{\bm{x}}}\sim\mathcal{P}_{\tilde{{\bm{x}}}}$). Then, the following holds,
[…]
@@
Recall that the models are trained using noisy data sampled from $\mathcal{P}_{{\bm{x}}}$ and tested on clean data drawn from $\mathcal{P}_{\tilde{{\bm{x}}}}$ to see if there is noise memorization (high test loss) or signal learning (low test loss). For simplicity of notation, we will omit the attributions $a$ and $b$ unless explicitly needed. Without the loss of generality, we denote the training set by ${\bm{X}}^{D\times N}=({\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{N})$ where the matrix column ${\bm{x}}_{i}^{{D\times 1}}$ are w.r.t formulation (1). Recall that $\text{rank}({\bm{X}})=D,$ with probability $1$ under assumption C.1. Let $\bm{\Sigma}\in\mathbb{R}^{D\times D}$ represent the covariance matrix, defined as $\bm{\Sigma}={\bm{X}}{\bm{X}}^{\top}$. Let its singular value decomposition (SVD) be expressed as $\bm{\Sigma}={\bm{U}}\bm{\Lambda}{\bm{U}}^{\top}$, and similarly $\bm{\Sigma}_{\tilde{{\bm{x}}}}^{{D\times D}}=\tilde{{\bm{U}}}\bm{\Omega}\tilde{{\bm{U}}}^{\top}$. Note that the eigenvectors of $\bm{\Sigma}$, which are the columns of $\bm{U}$}, are arranged in descending order corresponding to their associated eigenvalues in $\bm{\Lambda}$. By Lemma C.2, since $\text{rank}({\bm{X}})=D\geq m$, then for the optimal solution, $\bm{\theta}^{opt}=\text{argmin}_{\bm{\theta}}\mathcal{L}(\bm{\theta};{\bm{X}}),$ we have that $\operatorname{rank}\left({\bm{W}}^{opt}_{L-1}\cdots{\bm{W}}^{opt}_{2}\right)=m,$ means that the rank of the inner layers multiplication is exactly the size of the bottleneck. By Theorem 3.2, proof Case II in (Kawaguchi, 2016), all local minima of the considered optimization problem are global minima, and the expression of a global minimum of the training optimization, namely
$\bm{\theta}^{opt}=\text{argmin}_{\bm{\theta}}\mathcal{L}(\bm{\theta};{\bm{X}})=\text{argmin}_{\bm{\theta}}\frac{1}{N}\|\Phi({\bm{X}};\bm{\theta})-{\bm{X}}\|_{F}^{2},$
@@
$\bm{\theta}^{opt}={\bm{U}}_{m}{\bm{U}}_{m}^{\top}{\bm{X}}{\bm{X}}^{\top}\left({\bm{X}}{\bm{X}}^{\top}\right)^{-},$
where ${\bm{U}}_{m}^{{D\times m}},$ is the matrix containing the first $m$ eigenvectors of the full eigenbasis ${\bm{U}}^{{D\times D}}$ of ${\bm{X}}{\bm{X}}^{\top},$ namely ${\bm{U}}_{m}=\left[{\bm{u}}_{1},\ldots,{\bm{u}}_{m}\right]$ (those corresponding to the $m$ largest eigenvalues), and the operator $(\cdot)^{-}$ stands for the pesudo inverse. Since $\bm{\Sigma}^{{D\times D}}={\bm{X}}{\bm{X}}^{\top}$ is invertible ($\text{rank}({\bm{X}})=D$), we have
$\bm{\theta}^{opt}={\bm{U}}_{m}{\bm{U}}_{m}^{\top},$
@@
$\Phi({\bm{X}};\bm{\theta})={\bm{U}}_{m}{\bm{U}}_{m}^{\top}{\bm{X}}.$
Since both, ${\bm{U}}^{{D\times D}}$ and $\tilde{{\bm{U}}}^{{D\times D}}$ are orthonormal and span $\mathbb{R}^{D}$ (by the SVD property), there exists an orthogonal matrix ${{\bm{R}}}^{D\times D}$ such that ${\bm{U}}={{\bm{R}}}\tilde{{\bm{U}}}$. This matrix ${{\bm{R}}}={\bm{U}}\tilde{{\bm{U}}}^{\top}$ is unitary by construction, satisfying ${{\bm{R}}}^{\top}{{\bm{R}}}=\mathbb{I}_{D\times D}$, and represents a rotation or reflection aligning the test and training eigenspaces. Thus $\bm{U}$} and $\tilde{{\bm{U}}}$ are related by this $\bm{R}$}}, such that
${\bm{U}}={{\bm{R}}}\tilde{{\bm{U}}}$.
Although the truncated matrices ${\bm{U}}_{m}$ and $\tilde{{\bm{U}}}_{m}$ generally span different $m$-dimensional subspaces, both reside within the same ambient $D$-dimensional space, thus enabling such an orthonormal transformation $\bm{R}$}}.
@@
[…]
$=\mathbb{E}\left[\|{\bm{U}}_{m}{\bm{U}}_{m}^{\top}\tilde{{\bm{x}}}-\tilde{{\bm{x}}}\|_{2}^{2}\right]$
$=\mathbb{E}\left[\|\right(\mathbb{I}-{\bm{U}}_{m}{\bm{U}}_{m}^{\top}\left)\tilde{{\bm{x}}}\|_{2}^{2}\right]$
$={\mathbb{E}\left[\tilde{{\bm{x}}}^{T}(\mathbb{I}-{\bm{U}}_{m}{\bm{U}}_{m}^{T})^{T}(\mathbb{I}-{\bm{U}}_{m}{\bm{U}}_{m}^{T})\tilde{{\bm{x}}}\right]}$
$={\mathbb{E}\left[\text{Tr}\left((\mathbb{I}-{\bm{U}}_{m}{\bm{U}}_{m}^{T})\tilde{{\bm{x}}}\tilde{{\bm{x}}}^{\top}\right)\right]}$
$={\text{Tr}\left[(\mathbb{I}-{\bm{U}}_{m}{\bm{U}}_{m}^{T})\mathbb{E}\left[\tilde{{\bm{x}}}\tilde{{\bm{x}}}^{\top}\right]\right]}$
$={\text{Tr}\left[(\mathbb{I}-{\bm{U}}_{m}{\bm{U}}_{m}^{T})\bm{\Sigma}_{\tilde{{\bm{x}}}}\right]}$
$={\text{Tr}\left[\bm{\Sigma}_{\tilde{{\bm{x}}}}\right]-\text{Tr}\left[{\bm{U}}_{m}{\bm{U}}_{m}^{T}\bm{\Sigma}_{\tilde{{\bm{x}}}}\right]}$
$={\text{Tr}\left[\bm{\Sigma}_{\tilde{{\bm{x}}}}\right]-\text{Tr}\left[{{\bm{R}}}\tilde{{\bm{U}}}_{m}\tilde{{\bm{U}}}_{m}^{\top}{{\bm{R}}}^{\top}\bm{\Sigma}_{\tilde{{\bm{x}}}}\right]}$
$=\mathrm{Tr}(\bm{\Sigma}_{\tilde{{\bm{x}}}})-\mathbb{E}\left[\left\|\tilde{{\bm{U}}}_{m}^{\top}{{\bm{R}}}^{\top}\tilde{{\bm{x}}}\right\|_{2}^{2}\right],$ (4)
where $\bm{\Sigma}_{\tilde{{\bm{x}}}}^{{D\times D}}:=\mathbb{E}\left[\tilde{{\bm{x}}}\tilde{{\bm{x}}}^{\top}\right$]. Now, since there exists at least $d$ different indexes $j\in[D]$ for which ${\bm{u}}_{j}^{\top}{{\bm{R}}}\neq 0$ (otherwise the latent dimension of the test-set $\tilde{{\bm{x}}}$ would be strictly lower than $d$ in contradiction), then for $m_{a}+1<d$ and $m_{b}+1<d$, $\mathbb{E}\left[\left\|\tilde{{\bm{U}}}_{m_{a}}^{\top}{{\bm{R}}}^{\top}\tilde{{\bm{x}}}\right\|_{2}^{2}\right],\quad\mathbb{E}\left[\left\|\tilde{{\bm{U}}}_{m_{b}}^{\top}{{\bm{R}}}^{\top}\tilde{{\bm{x}}}\right\|_{2}^{2}\right]>0.$ Hance, since the larger is $m$ in (4) the lower is $\mathcal{G},$ we have
[…]
@@
$\mathcal{G}\left(\mathcal{D}_{b};\bm{\theta_{b}}^{\text{opt }},\mathcal{P}_{\tilde{{\bm{x}}}}\right)=\mathcal{G}\left(\mathcal{D}_{a};\bm{\theta_{a}}^{\text{opt }},\mathcal{P}_{\tilde{{\bm{x}}}}\right)=0.$
@@
Consider the pre-trained linear AE, $\Phi(\cdot;\bm{\theta}^{\text{opt}}):\mathbb{R}^{D}\to\mathbb{R}^{D}$, of $L\geq 2$ layers, where $\bm{\theta}^{\text{opt}}\equiv\left\{{\bm{W}}_{1}^{d_{1}\times D},\ldots,{\bm{W}}_{L-1}^{d_{L-1}\times d_{L-2}},{\bm{W}}_{L}^{D\times d_{L-1}}\right\}$, for which its bottleneck size holds $m(\leq D)$, and a training set $\bm{X}$}, of any dimension satisfies $\text{rank}\left({\bm{X}}\right)\geq m,$ then $\operatorname{rank}\left({\bm{W}}_{L-1}\cdots{\bm{W}}_{2}\right)=m.$